OpenStudy (anonymous):

Which of the following demonstrates the Commutative Property of Multiplication? 2(6a − 3) = (6a − 3) ⋅ 2 2(6a − 3) = 12a − 6 12a − 6 = (6a − 3) ⋅ 2 (2 ⋅ 6a) − 3 = 2(6a − 3)

OpenStudy (anonymous):

@AakashSudhakar @Miracrown @Elsa213 @Toff @ikram002p

OpenStudy (anonymous):

Fan and medal

pooja195 (pooja195):

\[\huge~\rm~a~\times~b=b~\times~a\]

OpenStudy (anonymous):

So c? @pooja195

pooja195 (pooja195):

I think thats right....im not sure :( @Nnesha ?

OpenStudy (irishboy123):

crumbs almighty. 2(6a − 3) = (6a − 3) ⋅ 2 is actually commuting the others are distributing or are wrong

pooja195 (pooja195):

Aight so C was correct :)

OpenStudy (anonymous):

Ok thank you i have to give the medal to pooja since he/she led me to the answer but ill fan both of you thanks :)

OpenStudy (irishboy123):

seriously: only a is commuting c is doing a bit more ...... and medals are for wimps!!!! have a good evening!!!

OpenStudy (anonymous):

thank you!!!!!!! you too!!!!!!!!!! :3